OpenStudy (anonymous):

Here's another system of equations: y = 2 - x -5x - 15y = 60 Which method would you use to solve this system? What is the solution to this system of equations?

OpenStudy (anonymous):

I would use substitution. substitute "2 - x" in for y in the second equation and solve for x. then sub that into the first equation to get y.

jhonyy9 (jhonyy9):

check it like : -5x -15(2-x))=60 -5x-30+15x=60 10x=90 x=9 y=2-9=-7 x=9 y=-7